A 20g mass is suspended from a metre rod at the 20cm.The metre rod is balanced on 40cm mark on wedge.Weight of metre rod is?
1
Expert's answer
2017-10-23T14:40:07-0400
The pivot point is at the 40 cm mark. The weight of the 20 gram object is at the 20 cm mark.
Counter clockwise torque:
0.02(10)(0.4-0.2) =0.04
The weight of the rod is at the center of the rod.
Clockwise torque:
W(0.5-0.4)=0.1W
The weight of meter rod is
W=0.04/0.1=0.4 N.